A certain quantity of raisins contains 308 mg. of potassium. An equal amount of cranberries contains 44 mg. of potassium. A 400-gram mixture of the two is made that ends up containing 220 mg. of potassium. How many grams of raisins were used in the mixture? How many grams of cranberries were used? This is a method of elimination problem but I can't seem to eliminate anything successfully from the two equations that I found. Please help.
Answer by jorel1380(3719)

If you're making a 400 gram mixture of cranberries and raisins, then C+R=400. Let's take one gram of raisins---It's made up of .308 grams of potassium. Then:
R+C=400gm
.308R+.044C=.220(400gm)=88gm
Multiply the second equation by 1000
308R+44C=88000
44R+44C=400X44=17600
(308-44)R=88000-17600=70400
264R=70400
R=266 2/3 gms.
C=133 1/3 gms.